# Proximity and Visuotactile Point Cloud Fusion for Contact Patches in Extreme Deformation

### v1

@@ VII-C Pose Estimation
In this demonstration, we show that our contact patches are compatible with a standard pose estimation pipeline. Because our contact patch algorithm provides accurate contact patches in high deformation of the membrane, the greater conformal contact allows for large and dense contact patch point clouds that are easy to use with a standard Iterative Closest Point (ICP) implementation from Open3D [50] to estimate pose. Exclusively using contact patches for pose estimation, rather than only the proximity modality, could be advantageous in cluttered environments where an additional segmentation step would be necessary to isolate the object of interest. We estimate the pose of a cup for 5 s as the cup simultaneously rotates about the handle and translates across the sensor surface. Because of stiction between the cup and the sensor surface, wrinkles are created in the membrane during the cup trajectory. We assume a known model of the cup geometry. In the first frame of the trajectory, the cup is manually registered to the estimated contact patch. For the following frames that measure the rest of the trajectory, the relative transformation of the cup is estimated for each timestep.
We observe that the key feature of the cup that leads to stable pose estimations is visibility of the handle in the contact patch. The pose estimation tends to become unstable when the handle is not observable and when there are fewer points in the contact patch estimation. Qualitatively, the pose estimate of the cup matches very well (Figure 10) and is fairly stable across the 5 s of data. Because the external environment was significantly brighter than the internal sensor chamber in this experiment, the tactile depth camera struggled to accurately estimate the membrane point cloud, leading to noise in the contact patch from the proximity depth sensing the hand holding the cup. The sensor could be modified to increase robustness to external light by increasing the intensity of the internal LEDs in the sensor chamber and formulating a more visibly-opaque membrane by adding more dye.
A future application of this work could be tactile SLAM for simultaneous object pose estimation and reconstruction [51, 52].
@@ VIII Conclusion and Future Work
In this study, we introduced a proximity and visuotactile point cloud fusion algorithm to detect contact patches with soft membranes across an unprecedented range and type of deformation. The key idea of the algorithm is that the contact patch can be identified as the set of points where the proximity point cloud and tactile point cloud intersect. Because the proximity point cloud provides localization and a point cloud of the object geometry independent of membrane mechanics, our method is not limited by the difficulties of accurately modeling the complex deformations in hyperelastic membrane materials that occur during real-world object interactions. This is not possible for visuotactile sensors that rely exclusively on the tactile modality.
We also presented the design and fabrication techniques for a selectively transmissive soft membrane that enables the dual sensing modalities. The formulation of the soft membrane leverages the insolubility of a visibly-colored, IR-transparent aqueous dye solution in a highly elastic silicone elastomer to create visual texture for tactile sensing with a stereo RGB depth camera and enable proximity sensing for an internal ToF depth camera. The fundamental sensing mechanism of selective light transmission is robust to manufacturing variances in the membrane, so the proposed sensor fusion algorithm requires no tuning for different batches of membranes. Additionally, the fabrication process uses accessible, off-the-shelf materials and produces a physically resilient and easily repairable sensing interface.
Furthermore, we conducted an extensive study of the accuracy of contact patch geometry for a diverse object dataset across three regimes of deformation: low strain, medium strain, and high strain. We found that the proposed proximity and visuotactile fusion algorithm used with the selectively transmissive membrane had the best performance across all three strain categories.
We demonstrated the use of this sensor and algorithm with three applications: use cases with complex membrane mechanics, closed loop control for balancing a stack of books with a dual arm manipulation platform, and pose estimation.
Moving forward, we would like to explore modifying the sensor design to achieve higher sampling frequencies for more dynamic manipulation tasks. The current sensor is limited by the frame rate of the proximity ToF depth camera, to maintain synchronous sampling with the tactile stereo RGB depth camera. The selective transmission sensing mechanism does not solely rely on using a ToF depth camera; the requirement is for an IR camera. A high-speed IR camera (stereo or mono) could be used to replace the ToF depth camera to increase the speed of the entire sensing system.
This work could also be extended by using this contact patch algorithm and selective light transmission sensing mechanism for variable stiffness surfaces. As the soft robotics field moves towards multifunctional materials with tighter integration of soft sensors and actuators, a natural next step for this work could be integration with an actuating surface that changes stiffness and shape in real-time to manipulate objects.
Another future direction for this work is to further investigate the use of proximity depth sensing in perception for manipulation. In this paper, we presented proximity as a modality to support and enhance tactile sensing functionality. It would be interesting to study the benefits of proximity sensing as a more independent modality in a manipulation pipeline that incorporates vision and tactile sensing.
@@ Acknowledgment
The authors would like to thank the PrOps team, Aimee Goncalves, Dr. Jose Barrieros, Dr. Aykut Önol, Eric Cousineau, and Alex Alspach at the Toyota Research Institute for helpful discussions, assistance with setting up experiments, and hardware design advice. The authors are grateful for assistance with running experiments and insightful discussions about the pose estimation demonstration from William Yang at the University of Pennsylvania. The authors also thank Gregory Campbell from the University of Pennsylvania for assistance with running experiments.

### v2

@@ VII-C Pose Estimation
In this demonstration, we show that our contact patches are compatible with a standard pose estimation pipeline (Fig. 7). The greater conformal contact allows for large and dense contact patch point clouds that are easy to use with a standard Iterative Closest Point (ICP) implementation from Open3D [32] to estimate pose. Exclusively using contact patches for pose estimation, rather than only the proximity modality, could be advantageous in cluttered environments where an additional segmentation step would be necessary to isolate the object of interest. We estimate the pose of a cup for 5 s as the cup simultaneously rotates about the handle and translates across the sensor surface. We assume a known model of the cup geometry. In the first frame of the trajectory, the cup is manually registered to the estimated contact patch. For the following frames that measure the rest of the trajectory, the relative transformation of the cup is estimated for each timestep. The pose estimation tends to become unstable when the cup handle is not observable. Qualitatively, the pose estimate of the cup matches well and is fairly stable across the 5 s of data. A future application of this work could be tactile SLAM for simultaneous object pose estimation and reconstruction [33, 34].
@@ VIII Conclusion and Future Work
In this study, we introduced a proximity and visuotactile point cloud fusion algorithm to detect contact patches with soft membranes across an unprecedented range and type of deformation. The key idea of the algorithm is that the contact patch can be identified as the set of points where the proximity point cloud and tactile point cloud intersect. Because the proximity point cloud provides localization and object geometry independent of membrane mechanics, our method is not limited by the challenge of accurately modeling the complex deformations in hyperelastic membrane materials.
We found that our proposed proximity and visuotactile fusion algorithm offered the best performance across all three strain categories. We demonstrated the use of this algorithm with three applications: complex membrane mechanics such as varied stiffness and wrinkled membranes, closed loop control for balancing a stack of books with a dual arm manipulation platform, and pose estimation.
One future direction is to explore a fingertip form factor. The current sensor size is driven by minimum sensing distances of the internal depth cameras, which can be replaced with shorter-range sensors for a significantly smaller overall package. Another future direction is to investigate using this sensor and algorithm with an actuating surface that changes stiffness and shape in real-time. A shape-changing end effector can be optimized for interactions with different objects. Our algorithm is uniquely suited to mitigate the challenges of using soft materials with variable stiffnesses and shapes.
@@ Acknowledgment
We thank the PrOps team, Aimee Goncalves, Dr. Jose Barrieros, Dr. Aykut Önol, Eric Cousineau, and Alex Alspach at TRI for helpful discussions, assistance with experiments, and hardware design advice. We thank William Yang and Greg Campbell at the Penn GRASP Lab for assistance with experiments and helpful discussions.